OpenStudy (lovelyanna):

Which relation is a function? A. {(1, 2); (1, 3); (3, 2); (4, 2)} B. {(1, 2); (2, 5); (3, 2); (4, 5)} C. {(1, 2); (2, 3); (3, 4); (2, 5)} D. {(1, 2); (2, 3); (3, 4); (1, 5)} B?

OpenStudy (anonymous):

Functions are for each value of x (the first number in an ordered pair) there can only be one y value (the second number in an ordered pair).

OpenStudy (jdoe0001):

a function will have NO X-REPEATS

OpenStudy (math&ing001):

Yeah, so B is correct

OpenStudy (lovelyanna):

Thanks! I cant afford to get any wrong! Hi @math&ing001

OpenStudy (lovelyanna):

How come I can't tag your name?

OpenStudy (math&ing001):

Hey there :)

OpenStudy (lovelyanna):

Sorry.. computer glitching.. how are you? Do you think you can help me with other questions? I don't want to get any wrong!

OpenStudy (lovelyanna):

@math&ing001 ?

OpenStudy (math&ing001):

Yeah, sure

OpenStudy (lovelyanna):

Thanks! For the function f(x) = 2 – 3x, find f(0). A. 2 B. 5 C. –1 D. –3 Its a bit confusing.. but I don't think the answer is positive?

OpenStudy (math&ing001):

You just replace the x with 0 like this: f(0) = 2 - 3*0

OpenStudy (lovelyanna):

Oh.. then its A? -.- Wow..

OpenStudy (math&ing001):

Yep ^^

OpenStudy (lovelyanna):

Thanks! :) Which is the equation of the function graphed? A. f(x) = |x| – 4 B. f(x) = |x – 4| C. f(x) = |x| + 4 D. f(x) = |x + 4|

OpenStudy (math&ing001):

Can you try to figure it out ? Plug in some values in your functions and see which one corresponds to the graph given.

OpenStudy (lovelyanna):

Umm.. I'm not soo good at graphs.. would it be x + 4.. I'm not sure where the lines go

OpenStudy (math&ing001):

Do you know what this sign means |x| ?

OpenStudy (lovelyanna):

Ya.. but I don't know how to explain.. like when |-9| its 9.. right?

OpenStudy (math&ing001):

Yeah it's the absolute value of x meaning the non-negative value of x. For exemple f(-2)=|-2|+4=2+4=6

OpenStudy (lovelyanna):

I see.. would it be D..?

OpenStudy (math&ing001):

Nope, but how did you get it ?

OpenStudy (lovelyanna):

I'm still confused.. and I don't know :/ but the answer is x + 4 which is either C or D?

OpenStudy (math&ing001):

Ok plug in 0 for both choices and see what it gives you

OpenStudy (lovelyanna):

You mean for f(x)?

OpenStudy (math&ing001):

yep

OpenStudy (lovelyanna):

Umm.. give me a minute

OpenStudy (lovelyanna):

My head hurts.. can you explain more, please?

OpenStudy (math&ing001):

I'm gonna do the first one, you do the second one. D - f(0) = | 0 + 4 | = |4| = 4

OpenStudy (lovelyanna):

f(0) = |0| + 4 = 4 same answer..? :/

OpenStudy (math&ing001):

Yeah, correct !And they both belong to the graph. So we try another point say -4. Let's see what we get now.

OpenStudy (lovelyanna):

You mean -4 for x?

OpenStudy (math&ing001):

Exactly

OpenStudy (lovelyanna):

for C - |-4| + 4 = 4 + 4 = 8 for D - |-4 + 4| = |0| = 0 correct?

OpenStudy (math&ing001):

Yeah ! So which function is graphed

OpenStudy (lovelyanna):

C...?

OpenStudy (math&ing001):

Yep !

OpenStudy (lovelyanna):

Yes, finally! thanks :) more graphs up ahead.. >,<

OpenStudy (lovelyanna):

Which is the graph of the function g(x) = –|x + 3|?

OpenStudy (lovelyanna):

what are we supposed to replace for x?

OpenStudy (math&ing001):

You can choose any value you want. Always start easy like 0 for example and narrow down your choices.

OpenStudy (lovelyanna):

ok, so from 0 and go down, negatives?

OpenStudy (lovelyanna):

can you give me one example?

OpenStudy (math&ing001):

It doesn't matter, those are just random value that help you identify your graph. Example: g(1)= - | 1+3 | = -4 Only graph 3 checks, so graph 3 is the one we're looking for.

OpenStudy (lovelyanna):

I see... I'm sorry, but I just not good at graphs :(

OpenStudy (lovelyanna):

Which equation represents a direct linear variation? A. y = 1/2x B. y = x + 2 C. 1/x D. y = x^2 B?

OpenStudy (math&ing001):

No worries :)

OpenStudy (math&ing001):

I gotta go sleep it's 12 am already :p

OpenStudy (lovelyanna):

Oh, ok, bye then! You know anyone else who can help me, like you do?

OpenStudy (math&ing001):

There's @hartnn but I'm not sure he's online atm

OpenStudy (lovelyanna):

Oh, ok.. can you answer the last question above? I said it was B :P

OpenStudy (math&ing001):

Yep, that's correct !

OpenStudy (lovelyanna):

Ok, thanks! I don't know anyone on who will help me like you do :P Bummer..

OpenStudy (math&ing001):

I know a bunch but they're not online right now, well good night :)

OpenStudy (lovelyanna):

Ok.. bye!